two trains leave stations 224 miles apart at the same time and travel toward each other. one train travels at 80 miles per hour while the other travels at 60 miles per hour. how long will it take for the two trains to meet?
Found 2 solutions by richwmiller, Maths68:
Answer by richwmiller(17219)

60t+80t=224
140t=224
t=224/140
t=1.6 hrs
t=1 3/5 hrs
t=1 hour 36 minutes

Answer by Maths68(1474)

Train A
=========
Speed = 80 mph
Distance = x
Time=t
Time = Distance / speed
Time = x/80
Train B
=========
Speed = 60mph
Distance = 224-x
Time = t
Time = Distance / speed
Time = 224-x/60

Both train will travel for the same period of time so
Time taken train A = Time taken train B
(224-x)/60=x/80
80(224-x)=60x
17920-80x=60x
17920=60x+80x
17920=140x
x=17920/140
x=128
Time taken by train A = x/80=128/80=1.6 hour
Time taken by train B = 224-128/80=96/60=1.6 hour
It will take 1.6 hour when they meet.


2nd Method
===========
Effective Speed = 80+60=140 mph
Distance = 224 miles
Time = Distance/Speed
Time = 224/140
Time = 1.6 hour